# Reinforced Lifelong Editing for Language Models

### v1

@@ Appendix B Related Work
Parameter-modifying Editing. Parameter-modifying methods modify LLM parameters through various approaches to achieve knowledge editing. One approach is the locate-then-edit method. Knowledge-neuron (Dai et al., 2022) pioneered the use of causal analysis to locate knowledge in LLMs. ROME (Meng et al., 2022) adopted this idea, using causal trace to identify factual associations and edit knowledge in MLP layers. MEMIT (Meng et al., 2023) extended this approach to mass-editing. PMET (Li et al., 2024) specifically investigated the role of MHSA in model editing. ECE (Zhang et al., 2025) incorporates LLM explainability into the editing process and clusters similar knowledge based on explanation results. Another approach is meta-learning. KE (De Cao et al., 2021) first proposed using hypernetwork for knowledge editing. MEND (Mitchell et al., 2022a) applied low-rank decomposition to LLM fine-tuning gradients, enabling faster and more effective training in hypernetworks. MALMEN (Tan et al., 2024) aggregated parameter shifts using normal equations, extending hypernetworks to mass-editing.
@@ C.2 More Results of Editing with Varying Numbers of Knowledge
Task CounterFact ZsRE FEVER Method Dataset Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ 20 ✕ 100 = 2000 FT 83.33$\pm$0.37 67.79$\pm$0.40 46.63$\pm$0.37 30.54$\pm$0.27 30.29$\pm$0.27 15.49$\pm$0.18 7.35$\pm$0.18 6.00$\pm$0.16 24.10$\pm$0.15 MEND 49.20$\pm$0.42 50.10$\pm$0.33 50.05$\pm$0.23 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 37.75$\pm$0.21 37.78$\pm$0.30 29.10$\pm$0.27 ROME 64.45$\pm$0.37 61.42$\pm$0.44 49.46$\pm$0.40 2.00$\pm$0.11 1.68$\pm$0.15 0.68$\pm$0.07 MEMIT 65.05$\pm$0.48 64.68$\pm$0.43 52.33$\pm$0.39 57.72$\pm$0.37 52.48$\pm$0.37 25.78$\pm$0.22 PRUNE 68.25$\pm$0.28 64.75$\pm$0.01 49.82$\pm$0.24 24.77$\pm$0.37 23.87$\pm$0.03 20.69$\pm$0.43 RECT 64.00$\pm$0.48 61.20$\pm$0.43 60.88$\pm$0.37 86.02$\pm$0.24 81.81$\pm$0.27 32.04$\pm$0.23 AlphaEdit 98.90$\pm$0.10 94.22$\pm$0.19 67.88$\pm$0.29 94.47$\pm$0.13 91.13$\pm$0.19 32.55$\pm$0.22 - RLEdit 91.75$\pm$0.21 62.40$\pm$0.18 52.38$\pm$0.19 88.65$\pm$0.19 83.91$\pm$0.24 47.61$\pm$0.23 94.46$\pm$0.18 91.56$\pm$0.29 69.01$\pm$0.29 50 ✕ 100 = 5000 FT 82.10$\pm$0.19 66.08$\pm$0.29 40.35$\pm$0.36 21.05$\pm$0.24 20.80$\pm$0.24 9.69$\pm$0.14 8.40$\pm$18.72 6.01$\pm$16.38 23.69$\pm$0.15 MEND 49.71$\pm$0.32 49.63$\pm$0.33 50.23$\pm$0.49 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 33.49$\pm$0.31 28.26$\pm$0.22 26.77$\pm$0.35 ROME 48.62$\pm$0.50 49.78$\pm$0.48 51.65$\pm$0.46 1.25$\pm$0.11 1.30$\pm$0.11 1.65$\pm$0.06 MEMIT 64.21$\pm$0.23 60.07$\pm$0.43 46.64$\pm$0.37 0.07$\pm$0.01 0.07$\pm$0.01 31.67$\pm$0.22 PRUNE 56.07$\pm$0.34 54.58$\pm$0.39 50.32$\pm$0.28 0.09$\pm$0.01 0.08$\pm$0.01 1.35$\pm$0.04 RECT 52.64$\pm$0.50 50.02$\pm$0.46 53.59$\pm$0.42 0.07$\pm$0.01 0.07$\pm$0.01 1.34$\pm$0.03 AlphaEdit 94.66$\pm$0.15 92.35$\pm$0.22 62.00$\pm$0.31 93.76$\pm$0.15 88.65$\pm$0.21 31.71$\pm$0.22 RLEdit 85.54$\pm$0.43 60.96$\pm$0.27 48.60$\pm$0.29 85.46$\pm$0.39 80.68$\pm$0.44 43.35$\pm$0.21 94.37$\pm$0.18 91.84$\pm$0.22 67.58$\pm$0.27 100 ✕ 100 = 10000 FT 79.99$\pm$0.20 62.02$\pm$0.29 38.86$\pm$0.36 14.80$\pm$0.22 14.50$\pm$0.22 5.22$\pm$0.10 23.02$\pm$0.28 8.38$\pm$0.19 22.16$\pm$0.16 MEND 47.24$\pm$0.32 47.19$\pm$0.23 53.10$\pm$0.35 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 27.24$\pm$0.27 28.93$\pm$0.32 27.56$\pm$0.34 ROME 46.43$\pm$0.39 40.99$\pm$0.31 46.68$\pm$0.23 1.87$\pm$0.12 1.84$\pm$0.12 1.65$\pm$0.06 MEMIT 47.53$\pm$0.50 43.59$\pm$0.49 51.30$\pm$0.48 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 PRUNE 48.29$\pm$0.23 48.01$\pm$0.37 50.91$\pm$0.43 0.09$\pm$0.01 0.06$\pm$0.02 1.26$\pm$0.09 RECT 49.95$\pm$0.29 50.11$\pm$0.32 46.27$\pm$0.25 0.08$\pm$0.01 0.08$\pm$0.01 1.36$\pm$0.04 AlphaEdit 80.17$\pm$0.21 75.34$\pm$0.45 52.78$\pm$0.39 89.60$\pm$0.17 86.75$\pm$0.23 30.96$\pm$0.22 RLEdit 80.69$\pm$0.23 62.31$\pm$0.27 45.16$\pm$0.31 86.45$\pm$0.32 83.07$\pm$0.19 41.96$\pm$0.34 95.06$\pm$0.21 92.26$\pm$0.18 69.19$\pm$0.30 150 ✕ 100 = 15000 FT 89.79$\pm$0.43 70.05$\pm$0.32 36.64$\pm$0.22 13.95$\pm$0.20 13.49$\pm$0.20 3.93$\pm$0.08 44.84$\pm$0.43 32.19$\pm$0.42 33.54$\pm$0.41 MEND 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 21.70$\pm$0.43 20.30$\pm$0.24 13.86$\pm$0.25 ROME 46.51$\pm$0.50 47.10$\pm$0.47 53.36$\pm$0.45 1.47$\pm$0.11 1.43$\pm$0.11 0.72$\pm$0.06 MEMIT 52.07$\pm$0.45 50.23$\pm$0.48 48.89$\pm$0.47 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 PRUNE 52.19$\pm$0.24 52.87$\pm$0.39 44.43$\pm$0.27 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 RECT 54.49$\pm$0.30 53.24$\pm$0.31 46.22$\pm$0.27 0.02$\pm$0.01 0.03$\pm$0.01 1.02$\pm$0.21 AlphaEdit 79.35$\pm$0.35 69.28$\pm$0.17 42.01$\pm$0.35 84.43$\pm$0.25 78.28$\pm$0.30 28.48$\pm$0.21 RLEdit 87.81$\pm$0.43 85.13$\pm$0.22 40.98$\pm$0.37 95.11$\pm$0.26 91.85$\pm$0.24 68.59$\pm$0.29 200 ✕ 100 = 20000 FT 86.23$\pm$0.49 69.77$\pm$0.31 38.94$\pm$0.33 13.21$\pm$0.30 11.11$\pm$0.27 4.24$\pm$0.15 55.89$\pm$0.37 50.07$\pm$0.52 49.79$\pm$0.30 MEND 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 11.08$\pm$0.27 8.01$\pm$0.11 15.38$\pm$0.53 ROME 44.43$\pm$0.39 46.21$\pm$0.27 50.77$\pm$0.33 1.56$\pm$0.21 1.34$\pm$0.19 1.04$\pm$0.08 MEMIT 49.98$\pm$0.29 46.55$\pm$0.30 45.23$\pm$0.33 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 PRUNE 50.21$\pm$0.22 47.43$\pm$0.32 47.11$\pm$0.22 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 RECT 43.32$\pm$0.29 42.21$\pm$0.37 42.01$\pm$0.20 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 AlphaEdit 74.22$\pm$0.57 63.94$\pm$0.46 39.12$\pm$0.45 77.74$\pm$0.42 70.01$\pm$0.34 27.96$\pm$0.28 RLEdit 91.38$\pm$0.43 89.93$\pm$0.29 41.98$\pm$0.30 94.03$\pm$0.37 90.67$\pm$0.42 68.71$\pm$0.41
@@ Appendix F More Discussion
In our experiments, we found that different choices of LLM editing layers significantly impact editing effectiveness. While locate-then-edit methods (Meng et al., 2023; Gu et al., 2024; Fang et al., 2024) typically choose to edit the first few layers of LLM due to their optimal editing performance, we discovered that RLEdit performs best when editing the middle and final layers of LLM. This insight motivates further research on lifelong editing from the perspective of sequential knowledge localization.

### v2

@@ Appendix B Related Work
Parameter-modifying Editing. Parameter-modifying methods modify LLM parameters through various approaches to achieve knowledge editing. One approach is the locate-then-edit method. Knowledge-neuron (Dai et al., 2022) pioneered the use of causal analysis to locate knowledge in LLMs. ROME (Meng et al., 2022) adopted this idea, using causal trace to identify factual associations and edit knowledge in MLP layers. MEMIT (Meng et al., 2023) extended this approach to mass-editing. PMET (Li et al., 2024) specifically investigated the role of MHSA in model editing. ECE (Zhang et al., 2025b) incorporates LLM explainability into the editing process and clusters similar knowledge based on explanation results. Another approach is meta-learning. KE (De Cao et al., 2021) first proposed using hypernetwork for knowledge editing. MEND (Mitchell et al., 2022a) applied low-rank decomposition to LLM fine-tuning gradients, enabling faster and more effective training in hypernetworks. MALMEN (Tan et al., 2024) aggregated parameter shifts using normal equations, extending hypernetworks to mass-editing.
@@ C.2 More Results of Editing with Varying Numbers of Knowledge
Task CounterFact ZsRE FEVER Method Dataset Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ 20 ✕ 100 = 2000 FT 83.33$\pm$0.37 67.79$\pm$0.40 46.63$\pm$0.37 30.54$\pm$0.27 30.29$\pm$0.27 15.49$\pm$0.18 7.35$\pm$0.18 6.00$\pm$0.16 24.10$\pm$0.15 MEND 49.20$\pm$0.42 50.10$\pm$0.33 50.05$\pm$0.23 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 37.75$\pm$0.21 37.78$\pm$0.30 29.10$\pm$0.27 ROME 64.45$\pm$0.37 61.42$\pm$0.44 49.46$\pm$0.40 2.00$\pm$0.11 1.68$\pm$0.15 0.68$\pm$0.07 MEMIT 65.05$\pm$0.48 64.68$\pm$0.43 52.33$\pm$0.39 57.72$\pm$0.37 52.48$\pm$0.37 25.78$\pm$0.22 PRUNE 68.25$\pm$0.28 64.75$\pm$0.01 49.82$\pm$0.24 24.77$\pm$0.37 23.87$\pm$0.03 20.69$\pm$0.43 RECT 64.00$\pm$0.48 61.20$\pm$0.43 60.88$\pm$0.37 86.02$\pm$0.24 81.81$\pm$0.27 32.04$\pm$0.23 AlphaEdit 98.90$\pm$0.10 94.22$\pm$0.19 67.88$\pm$0.29 94.47$\pm$0.13 91.13$\pm$0.19 32.55$\pm$0.22 RLEdit 91.75$\pm$0.21 62.40$\pm$0.18 52.38$\pm$0.19 88.65$\pm$0.19 83.91$\pm$0.24 47.61$\pm$0.23 94.46$\pm$0.18 91.56$\pm$0.29 69.01$\pm$0.29 50 ✕ 100 = 5000 FT 82.10$\pm$0.19 66.08$\pm$0.29 40.35$\pm$0.36 21.05$\pm$0.24 20.80$\pm$0.24 9.69$\pm$0.14 8.40$\pm$18.72 6.01$\pm$16.38 23.69$\pm$0.15 MEND 49.71$\pm$0.32 49.63$\pm$0.33 50.23$\pm$0.49 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 33.49$\pm$0.31 28.26$\pm$0.22 26.77$\pm$0.35 ROME 48.62$\pm$0.50 49.78$\pm$0.48 51.65$\pm$0.46 1.25$\pm$0.11 1.30$\pm$0.11 1.65$\pm$0.06 MEMIT 64.21$\pm$0.23 60.07$\pm$0.43 46.64$\pm$0.37 0.07$\pm$0.01 0.07$\pm$0.01 31.67$\pm$0.22 PRUNE 56.07$\pm$0.34 54.58$\pm$0.39 50.32$\pm$0.28 0.09$\pm$0.01 0.08$\pm$0.01 1.35$\pm$0.04 RECT 52.64$\pm$0.50 50.02$\pm$0.46 53.59$\pm$0.42 0.07$\pm$0.01 0.07$\pm$0.01 1.34$\pm$0.03 AlphaEdit 94.66$\pm$0.15 92.35$\pm$0.22 62.00$\pm$0.31 93.76$\pm$0.15 88.65$\pm$0.21 31.71$\pm$0.22 RLEdit 85.54$\pm$0.43 60.96$\pm$0.27 48.60$\pm$0.29 85.46$\pm$0.39 80.68$\pm$0.44 43.35$\pm$0.21 94.37$\pm$0.18 91.84$\pm$0.22 67.58$\pm$0.27 100 ✕ 100 = 10000 FT 79.99$\pm$0.20 62.02$\pm$0.29 38.86$\pm$0.36 14.80$\pm$0.22 14.50$\pm$0.22 5.22$\pm$0.10 23.02$\pm$0.28 8.38$\pm$0.19 22.16$\pm$0.16 MEND 47.24$\pm$0.32 47.19$\pm$0.23 53.10$\pm$0.35 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 27.24$\pm$0.27 28.93$\pm$0.32 27.56$\pm$0.34 ROME 46.43$\pm$0.39 40.99$\pm$0.31 46.68$\pm$0.23 1.87$\pm$0.12 1.84$\pm$0.12 1.65$\pm$0.06 MEMIT 47.53$\pm$0.50 43.59$\pm$0.49 51.30$\pm$0.48 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 PRUNE 48.29$\pm$0.23 48.01$\pm$0.37 50.91$\pm$0.43 0.09$\pm$0.01 0.06$\pm$0.02 1.26$\pm$0.09 RECT 49.95$\pm$0.29 50.11$\pm$0.32 46.27$\pm$0.25 0.08$\pm$0.01 0.08$\pm$0.01 1.36$\pm$0.04 AlphaEdit 80.17$\pm$0.21 75.34$\pm$0.45 52.78$\pm$0.39 89.60$\pm$0.17 86.75$\pm$0.23 30.96$\pm$0.22 RLEdit 80.69$\pm$0.23 62.31$\pm$0.27 45.16$\pm$0.31 86.45$\pm$0.32 83.07$\pm$0.19 41.96$\pm$0.34 95.06$\pm$0.21 92.26$\pm$0.18 69.19$\pm$0.30 150 ✕ 100 = 15000 FT 89.79$\pm$0.43 70.05$\pm$0.32 36.64$\pm$0.22 13.95$\pm$0.20 13.49$\pm$0.20 3.93$\pm$0.08 44.84$\pm$0.43 32.19$\pm$0.42 33.54$\pm$0.41 MEND 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 21.70$\pm$0.43 20.30$\pm$0.24 13.86$\pm$0.25 ROME 46.51$\pm$0.50 47.10$\pm$0.47 53.36$\pm$0.45 1.47$\pm$0.11 1.43$\pm$0.11 0.72$\pm$0.06 MEMIT 52.07$\pm$0.45 50.23$\pm$0.48 48.89$\pm$0.47 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 PRUNE 52.19$\pm$0.24 52.87$\pm$0.39 44.43$\pm$0.27 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 RECT 54.49$\pm$0.30 53.24$\pm$0.31 46.22$\pm$0.27 0.02$\pm$0.01 0.03$\pm$0.01 1.02$\pm$0.21 AlphaEdit 79.35$\pm$0.35 69.28$\pm$0.17 42.01$\pm$0.35 84.43$\pm$0.25 78.28$\pm$0.30 28.48$\pm$0.21 RLEdit 87.81$\pm$0.43 85.13$\pm$0.22 40.98$\pm$0.37 95.11$\pm$0.26 91.85$\pm$0.24 68.59$\pm$0.29 200 ✕ 100 = 20000 FT 86.23$\pm$0.49 69.77$\pm$0.31 38.94$\pm$0.33 13.21$\pm$0.30 11.11$\pm$0.27 4.24$\pm$0.15 55.89$\pm$0.37 50.07$\pm$0.52 49.79$\pm$0.30 MEND 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 11.08$\pm$0.27 8.01$\pm$0.11 15.38$\pm$0.53 ROME 44.43$\pm$0.39 46.21$\pm$0.27 50.77$\pm$0.33 1.56$\pm$0.21 1.34$\pm$0.19 1.04$\pm$0.08 MEMIT 49.98$\pm$0.29 46.55$\pm$0.30 45.23$\pm$0.33 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 PRUNE 50.21$\pm$0.22 47.43$\pm$0.32 47.11$\pm$0.22 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 RECT 43.32$\pm$0.29 42.21$\pm$0.37 42.01$\pm$0.20 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 AlphaEdit 74.22$\pm$0.57 63.94$\pm$0.46 39.12$\pm$0.45 77.74$\pm$0.42 70.01$\pm$0.34 27.96$\pm$0.28 RLEdit 91.38$\pm$0.43 89.93$\pm$0.29 41.98$\pm$0.30 94.03$\pm$0.37 90.67$\pm$0.42 68.71$\pm$0.41
@@ Appendix F More Discussion
In our experiments, we found that different choices of LLM editing layers significantly impact editing effectiveness. While locate-then-edit methods (Meng et al., 2023) typically choose to edit the first few layers of LLM due to their optimal editing performance, we discovered that RLEdit performs best when editing the middle and final layers of LLM. We believe this insight motivates further research on lifelong editing from the perspective of sequential knowledge localization. Additionally, we will investigate the use of advanced approaches of LLM interpretability (He et al., 2024; Zhou et al., 2024) to identify and extract crucial layers within the model, enabling targeted editing and refinement for improved performance and adaptability.
Furthermore, in the future, we aim to apply the reinforcement learning paradigm to a broader range of model editing methods, such as locate-and-edit (Jiang et al., 2025) and memory-based (Wang et al., 2024) approaches. Additionally, to address the challenges of updating multi-hop and related knowledge, we plan to enhance RLEdit by integrating external knowledge bases or knowledge graphs (Chen et al., 2024; Zhang et al., 2025a; Chen et al., 2023), which is crucial for further advancements.